# WritingBench: A Comprehensive Benchmark for Generative Writing

### v3

@@ B.2 Guidance Pool
Table 10: Query refinement guidance pool introduced in Section B.5.
B.3 Query Refine Prompt
@@
Table 11: Query refinement prompt introduced in Section 3.1.1.
B.4 Criteria Generation Prompt
System Evaluation Prompt You are an expert evaluator with extensive experience in evaluating the response of a given query.
Table 12: Evaluation system prompt introduced in Section 3.2.
@@ B.4 Criteria Generation Prompt
    "1-2": "Low score description: Clearly deficient in this aspect, with significant issues.",
    "3-4": "Below average score description: Lacking in several important areas, with noticeable problems.",
    "5-6": "Average score description: Adequate but not exemplary, meets basic expectations with some minor issues.",
    "7-8": "Above average score description: Generally strong but with minor shortcomings.",
    "9-10": "High score description: Outstanding in this aspect, with no noticeable issues." },
Table 13: Criteria generation prompt introduced in Section 3.2.
B.5 Rubric-based Scoring Prompt
Scoring Prompt Evaluate the Response based on the Query and criteria provided.
@@ B.5 Rubric-based Scoring Prompt
Provide your evaluation based on the criteria:
@@
Provide reasons for each score, indicating where and why any strengths or deficiencies occur within the Response. Reference specific passages or elements from the text to support your justification.
Ensure that each reason is concrete, with explicit references to the text that align with the criteria requirements.
Scoring Range: Assign an integer score between 1 to 10
Output format:
Return the results in the following JSON format, Only output this JSON format and nothing else:
@@
Table 14: Scoring prompt introduced in Section 3.2.
Appendix C Experiment Results
C.1 Ablation of CoT in Creative content
To validate the impact of CoT reasoning on creative content generation, we conduct knowledge distillation experiments using DeepSeek-R1. Two variants of Qwen-2.5-32B-Instruct are developed: 1) Base model trained with CoT-formatted instructions, 2) ablated version (-w/o CoT) using direct response generation. These models are evaluated on the Literature & Art (D5) subset of WritingBench and EQBench, a specialized benchmark for fiction evaluation, to thoroughly examine their capabilities in generating creative content.
Models WritingBench-D4 EQBench
Deepseek-R1 8.55 84.99
Qwen-2.5-32B-Instruct 7.34 48.17
Qwen-2.5-32B-CoT 8.66 82.48
-w/o CoT 8.49 79.43
Table 15: Ablation of CoT in creative content on two benchmarks.
As shown in Table 15, the CoT-enhanced model achieves 8.66 (vs. 8.49) on WritingBench-D5 and 82.48 (vs. 79.43) on EQBench, showing CoT’s effectiveness in generating creative content.
C.2 Case Study
Figure 6: Scores of different models across various input lengths on the WritingBench.
Figure 7: Scores of different models across various output lengths on the WritingBench.

### v4

@@ C.3 Guidance Pool
C.4 Query Refine Prompt
Introduced in Section 3.1.1.
@@ C.4 Query Refine Prompt
C.5 Criteria Generation Prompt
Introduced in Section 3.2.
Evaluation System Prompt You are an expert evaluator with extensive experience in evaluating the response of a given query.
@@ C.5 Criteria Generation Prompt
   "1-2": "Low score description: Critical deficiencies and major issues that prevent adequate functionality.",
   "3-4": "Below average score description: Lacking with noticeable shortcomings that impact overall effectiveness and require improvement.",
   "5-6": "Average score description: Adequate but not exemplary, Baseline performance that meets essential requirements. Most models may achieve this score.",
   "7-8": "Above average score description: Strong performance characterized by competent execution, though minor refinements are needed to achieve excellence.",
   "9-10": "High score description: Exceptional performance with all aspects optimally addressed, demonstrating superior effectiveness and quality without any flaws."
C.6 Rubric-based Scoring Prompt
Introduced in Section 3.2.
Evaluation System Prompt You are an expert evaluator with extensive experience in evaluating the response of a given query.
Since the query and response may be very long, the {criteria} will appear twice: once before and once after the query and response.
Scoring Prompt Evaluate the Response based on the Query and Criteria provided following the Scoring Rules.
** Scoring Rules **
"1-2": "Low score description: Critical deficiencies and major issues that prevent adequate functionality."
"3-4": "Below average score description: Lacking with noticeable shortcomings that impact overall effectiveness and require improvement."
"5-6": "Average score description: Adequate but not exemplary, Baseline performance that meets essential requirements. Most models may achieve this score."
"7-8": "Above average score description: Strong performance characterized by competent execution, though minor refinements are needed to achieve excellence."
"9-10": "High score description: Exceptional performance with all aspects optimally addressed, demonstrating superior effectiveness and quality without any flaws."
-Provide reasons for each score by indicating specific strengths or deficiencies within the Response. Reference exact text passages to justify the score, ensuring that each reason is concrete and aligns with the criteria requirements while highlighting key gaps from the ideal answer.
-Be very STRICT and do not be misled by format or length; ensure that the Response is thoroughly evaluated beyond superficial appearances.
-Carefully discern whether the content of the Response is an illusion, appearing substantial but actually entirely fabricated.
-Sometimes the model may only provide an introduction or an overview without truly completing the query, which should be considered a failed response. Carefully discern this.
-Scoring Range: Assign an integer score between 1 to 10
** Output format **
(Remove symbols that interfere with JSON parsing, don’t use " inside reason)
Return the results in the following JSON format, Only output the following JSON format and nothing else:
  "score": an integer score between 1 to 10,
@@ C.6 Rubric-based Scoring Prompt
Provide your evaluation based on the criteria restated below:
@@
** Output format **
(Remove symbols that interfere with JSON parsing, don’t use " inside reason)
Return the results in the following JSON format, Only output the following JSON format and nothing else:
@@
Appendix D Limitations
Appendix E Impact
The introduction of WritingBench and its associated evaluation framework presents significant impacts across multiple dimensions of LLM research and application development:
  • Comprehensive AI Writing Evaluation Ecosystem WritingBench’s extensive query set spans a wide range of domains and requirements, making it an ideal resource for evaluating both general-purpose writing skills and domain-specific expertise. Researchers can use the benchmark to assess a model’s overall versatility or focus on specific domains, such as storytelling, scientific writing, or business communication. This dual capability allows for nuanced evaluations that cater to both academic research and practical applications, bridging the gap between theoretical advancements and real-world needs. By publicly releasing WritingBench, including its evaluation protocols, criteria generation tools, and critic models, we contribute to greater transparency and reproducibility in the field of LLM research. Researchers can replicate experiments, validate findings, and build upon our work to further refine evaluation methodologies. Please note that although every released query is manually reviewed to remove harmful content, the responses generated via API calls depend on the underlying model’s behavior, cannot be guaranteed to be fully safe or accurate, and should not be used for commercial purposes.
  • Facilitating Domain-Specific Research and Development With its inclusion of 100 subdomains, WritingBench provides a unique opportunity for domain-specific optimization. For example, industries such as legal writing, medical documentation, or marketing can evaluate model performance on relevant subsets to select the most suitable models for their needs. Furthermore, WritingBench’s query construction strategy can be leveraged to generate diverse queries when specialized evaluation datasets are lacking in a particular field. This capability supports targeted research the development of highly specialized writing tools. We welcome researchers and experts from all fields to join the discussion and collaborate on the development of this benchmark.
  • Advancing Reinforcement Learning and Adaptive Evaluation Frameworks The query-dependent evaluation framework introduced in WritingBench opens new avenues for integrating adaptive evaluation methods into reinforcement learning pipelines. The ability to generate instance-specific criteria dynamically allows for precise scoring mechanisms that can be used to train models through reward-based optimization. Additionally, the framework’s capacity to produce pairwise preference data by comparing predicts against adaptive criteria for improving model alignment with human preferences. This approach enhances the quality of generated text and accelerates the development of models capable of handling complex, multi-dimensional writing tasks.